मराठी
महाराष्ट्र राज्य शिक्षण मंडळएचएससी विज्ञान (सामान्य) इयत्ता ११ वी

Solve the following: Find the probability that a year selected will have 53 Wednesdays

Advertisements
Advertisements

प्रश्न

Solve the following:

Find the probability that a year selected will have 53 Wednesdays

बेरीज
Advertisements

उत्तर

Let A ≡ the event that a year selected has 53 Wednesdays

L ≡ the event that leap year is selected

N ≡ the event that non-leap year is selected

The required event will happen if any one of L ∩ A and N ∩ A occurs.

These events are mutually exclusive

∴ the required probability

= P(L ∩ A) + P(N ∩ A)

= `"P"("L")*"P"("A"/"L") + "P"("N")*"P"("A"/"N")`  ...(1)

There is one leap year in 4 consecutive years

∴ P(L) = `1/4`

`"P"("A"/"L")` = Probability that year has 53 Wednesdays given that it is a leap year

Leap year has 366 day

366 = 7 x 52 + 2

∴ there are 52 full weeks and 2 days.

These days can be Sunday, Monday; Monday, Tuesday; Tuesday, Wednesday; Wednesday, Thursday; Thursday, Friday; Friday, Saturday; Saturday, Sunday.

There are 7 possibilities and favourable cases are 2

∴ `"P"("A"/"L") = 2/7`

Since there are 3 non-leap years in 4 consecutive years,

P(N) = `3/4`

`"P"("A"/"N")` = Probability that year has 53 Wednesdays given that it is a non-leap year

Non-leap year has 365 days

365 = 7 x 52 + 1

∴ there are 52 full weeks and 1 day.

This day could be any day of the week days i.e., any one of 7 days.

The number of favourable case is 1.

∴ `"P"("A"/"N") = 1/7`

∴ from (1), the required probability = `1/4*2/7 + 3/4*1/7`

= `5/28`.

shaalaa.com
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 9: Probability - Miscellaneous Exercise 9 [पृष्ठ २१४]

APPEARS IN

बालभारती Mathematics and Statistics (Arts and Science) Part 1 [English] Standard 11 Maharashtra State Board
पाठ 9 Probability
Miscellaneous Exercise 9 | Q II. (13) | पृष्ठ २१४

व्हिडिओ ट्यूटोरियलVIEW ALL [2]

संबंधित प्रश्‍न

A die, whose faces are marked 1, 2, 3 in red and 4, 5, 6 in green is tossed. Let A be the event "number obtained is even" and B be the event "number obtained is red". Find if A and B are independent events.


Events A and B are such that `P(A) = 1/2, P(B) = 7/12 and P("not A or not B") = 1/4` . State whether A and B are independent?


The probabilities of solving a specific problem independently by A and B are `1/3` and `1/5` respectively. If both try to solve the problem independently, find the probability that the problem is solved.


A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, exactly two students solve the problem?


The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the chance that the problem will be solved, if they try independently?


Two dice are thrown together. Let A be the event 'getting 6 on the first die' and B be the event 'getting 2 on the second die'. Are the events A and B independent?


A bag contains 3 yellow and 5 brown balls. Another bag contains 4 yellow and 6 brown balls. If one ball is drawn from each bag, what is the probability that, both the balls are of the same color?


Solve the following:

If P(A ∩ B) = `1/2`, P(B ∩ C) = `1/3`, P(C ∩ A) = `1/6` then find P(A), P(B) and P(C), If A,B,C are independent events.


Solve the following:

Let A and B be independent events with P(A) = `1/4`, and P(A ∪ B) = 2P(B) – P(A). Find P(B)


Solve the following:

For three events A, B and C, we know that A and C are independent, B and C are independent, A and B are disjoint, P(A ∪ C) = `2/3`, P(B ∪ C) = `3/4`, P(A ∪ B ∪ C) = `11/12`. Find P(A), P(B) and P(C)


Solve the following:

A machine produces parts that are either good (90%), slightly defective (2%), or obviously defective (8%). Produced parts get passed through an automatic inspection machine, which is able to detect any part that is obviously defective and discard it. What is the quality of the parts that make it throught the inspection machine and get shipped?


10% of the bulbs produced in a factory are of red colour and 2% are red and defective. If one bulb is picked up at random, determine the probability of its being defective if it is red.


If A and B are independent events such that 0 < P(A) < 1 and 0 < P(B) < 1, then which of the following is not correct?


For a loaded die, the probabilities of outcomes are given as under:
P(1) = P(2) = 0.2, P(3) = P(5) = P(6) = 0.1 and P(4) = 0.3. The die is thrown two times. Let A and B be the events, ‘same number each time’, and ‘a total score is 10 or more’, respectively. Determine whether or not A and B are independent.


Refer to Question 1 above. If the die were fair, determine whether or not the events A and B are independent.


Two dice are thrown together and the total score is noted. The events E, F and G are ‘a total of 4’, ‘a total of 9 or more’, and ‘a total divisible by 5’, respectively. Calculate P(E), P(F) and P(G) and decide which pairs of events, if any, are independent.


A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`. Find: `"P"("A'"/"B'")`


Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: P1P2 


A and B are events such that P(A) = 0.4, P(B) = 0.3 and P(A ∪ B) = 0.5. Then P(B′ ∩ A) equals ______.


In Question 64 above, P(B|A′) is equal to ______.


If the events A and B are independent, then P(A ∩ B) is equal to ______.


Two independent events are always mutually exclusive.


If A and B′ are independent events, then P(A' ∪ B) = 1 – P (A) P(B')


Let A and B be two events. If P(A | B) = P(A), then A is ______ of B.


One card is drawn at random from a well-shuffled deck of 52 cards. In which of the following case is the events E and F independent?

E : ‘the card drawn is black’

F : ‘the card drawn is a king’


If P(A) = `3/5` and P(B) = `1/5`, find P(A ∩ B), If A and B are independent events.


Let A and B be independent events P(A) = 0.3 and P(B) = 0.4. Find P(A ∩ B)


Two events E and F are independent when which conditional probability condition holds, provided \[P(E) \ne 0\]?


For \[E=\{3,6\}\] when a die is thrown once, what is \[P(E)\]?


For \[F=\{2,4,6\}\] when a die is thrown once, what is \[P(F)\]?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×